# YOLO-MARL: You Only LLM Once for Multi-agent Reinforcement Learning

### v1

@@ Abstract
Advancements in deep multi-agent reinforcement learning (MARL) have positioned it as a promising approach for decision-making in cooperative games. However, it still remains challenging for MARL agents to learn cooperative strategies for some game environments. Recently, large language models (LLMs) have demonstrated emergent reasoning capabilities, making them promising candidates for enhancing coordination among the agents. However, due to the model size of LLMs, it can be expensive to frequently infer LLMs for actions that agents can take. In this work, we propose You Only LLM Once for MARL (YOLO-MARL), a novel framework that leverages the high-level task planning capabilities of LLMs to improve the policy learning process of multi-agents in cooperative games. Notably, for each game environment, YOLO-MARL only requires one time interaction with LLMs in the proposed strategy generation, state interpretation and planning function generation modules, before the MARL policy training process. This avoids the ongoing costs and computational time associated with frequent LLMs API calls during training. Moreover, the trained decentralized normal-sized neural network-based policies operate independently of the LLM. We evaluate our method across three different environments and demonstrate that YOLO-MARL outperforms traditional MARL algorithms.
11footnotetext: † These authors contributed equally to this work.
1 Introduction
@@
Multi-agent reinforcement learning (MARL) algorithms have proven to be a powerful framework for addressing complex decision-making problems in multi-agent systems. With the rising applications of multi-agent systems, such as mobile robots in warehouses and games requiring complex reasoning and strategy, it is increasingly crucial for individual agents to learn, cooperate, or compete in dynamic environments without a centralized decision-maker (Papoudakis & Schäfer, 2021). In cooperative Markov games, agents are trained to coordinate their actions to maximize the joint rewards. However, existing MARL algorithms face challenges in learning distributed policies for cooperative games. Moreover, they struggle with tasks characterized by sparse rewards, dynamic environment, and large action spaces, which can hinder efficient learning and agent collaboration.
LLMs have excelled as high-level semantic planners due to its in-context learning abilities and prior knowledge (Ahn et al., 2022). Zhang et al. (2023) and Kannan et al. (2024) directly use LLMs as embodied agents, which demonstrate LLMs’ planning ability in multi-robot system. There are also works concentrating on utilizing the LLMs to guide the reinforcement learning (RL) training to reach better performances. ELLM (Du et al., 2023) leverage LLMs to suggest a goal to assist RL training whereas Kwon et al. (2023) focusing on the alignment between the action provided by LLM and the RL policy. While these approaches show exciting potential for integrating LLM within policy training, they have yet to extend their methods on multi-agent scenarios. More importantly, utilizing LLMs as agents or integrating them into the RL training loop presents certain challenges. Repeated interactions with LLMs in long-episode tasks or complex environments—especially when using advanced LLMs like Claude-3.5 or GPT-o1 can be time-consuming and costly; it becomes intractable for tasks requiring training over tens of millions of steps. Additionally, there is a risk of intermittent disconnections with the LLM, which could disrupt the training process and affect the system’s stability.
@@
Built on the identified insights and challenges, we introduce YOLO-MARL, as shown in Fig. 1, an innovative approach that leverages the planning capabilities of LLMs to enhance MARL policy training. In particular, the major strength of our framework is that it requires only a one-time interaction with the LLM for each game environment. After the strategy generation, state interpretation and planning function generation modules, there is no need for further LLMs interaction during the MARL training process, which significantly reduces the communication and computational overhead of LLM inferences. Moreover, YOLO-MARL demonstrates its strong generalization capabilities and simplicity for application: with the proposed strategy generation and state interpretation modules, our approach is compatible with various MARL algorithms such as Yu et al. (2022), Rashid et al. (2018), Lowe et al. (2020), and requires only basic background understanding of a new game environment from the users. We also evaluate our framework in a sparser reward multi-agent environment: Level-Based Foraging environment (Papoudakis & Schäfer, 2021), and a highly strategic task environment: the StarCraft Multi-Agent Challenge environment (Samvelyan et al., 2019), together with the MPE environment (Lowe et al., 2020), and show that YOLO-MARL outperforms several MARL baselines. We also provide several ablation study results to demonstrate the function of each module in the proposed framework. To the best of our knowledge, YOLO-MARL is among one of the first trials that incorporates the high-level reasoning and planning abilities of LLMs with MARL, since very limited literature of LLM for MARL has been introduced so far (Sun et al., 2024).
@@
  • This framework synergizes the planning capabilities of LLMs with MARL to enhance the policy learning performance in challenging cooperative game environments. In particular, our approach exploits the LLM’s wide-ranging reasoning ability to generate high-level assignment planning functions to facilitate agents in coordination.
@@
An overview of YOLO-MARL is presented in Figure 1. All prompts, environments, and generated planning functions can be found in Appendix.
@@
Figure 1: Depiction of our framework YOLO-MARL. (a). Strategy Generation: We pass basic environment and task description into the LLM to get generated strategies for this specific environment. (b). State Interpretation: We process the global states so that the format of global states will be more structured and organized for better comprehension by the LLM. (c). Planning Function Generation: We chain together the environment and task description, LLM generated strategies and state interpretation function. These prompts are then fed into the LLM to generate a planning function for this environment. (d). MARL Training: The state interpretation function and the generated planning function are integrated into the MARL training process. The LLM is no longer required for further interaction after the Planning Function Generation.
2 Related Work
2.1 Multi-Agent Reinforcement Learning
MARL has gained increasing attention due to its potential in solving complex, decentralized problems. Centralized training with decentralized execution has become a popular framework for overcoming the limitations of independent learning. Methods like QMIX (Rashid et al., 2018) and MADDPG (Lowe et al., 2020) use centralized critics or value functions during training to coordinate agents, while allowing them to execute independently during testing. In cooperative environments, algorithms like COMA (Foerster et al., 2017) and VDN (Sunehag et al., 2017) enable agents to share rewards and act in a coordinated fashion to maximize joint rewards.Wang et al. (2024) introduce a new approach using language constraint prediction to tackle the challenge of safe MARL in the context of natural language. However, the existing MARL algorithms may not perform well in sparse reward environments and still struggle in learning fully cooperative policy in some environments. So far, only very limited literature of using LLM for MARL has been proposed (Sun et al., 2024), and it remains unclear whether and how can LLM be leveraged for MARL-based decision-making.
2.2 Large Language Models for Single-Agent RL and Decision-Making
Many existing works utilize LLMs as parts of RL training process. Du et al. (2023) enhance agents’ exploration by computing the similarity between suggested goals from LLMs and agents’ demonstrated behaviors. Carta et al. (2023) leveraging language-based goals from LLMs by generating actions conditioned on prompts during online RL. Kwon et al. (2023) provides scalar rewards based on suggestions from LLMs to guide RL training. However, most of these approaches haven’t explored their works in the context of Markov games and require extensive interactions with LLMs during training.
Gupta et al. (2022) utilize CLIP’s visual embedding to an agent exploring of environment. Fan et al. (2022) studies a multi-task RL problem, where an agent is tasked with completing MINEDOJO tasks. Ahn et al. (2022) proposes SayCan which grounds LLMs via value functions of pretrained skills to execute abstract commands on robots. Liang et al. (2023) finds that code-writing LLMs can be re-purposed to write robot policy code. Huang et al. (2022) shows that by leveraging environment feedback, LLMs are able to form an inner monologue that allows them to more richly process and plan. Other research such as Ma et al. (2024) and Xie et al. (2023) use LLMs prior knowledge and code generation capability to generate reward functions, whereas we utilize code generation for planning functions. Lin et al. (2024) highlights the limitations of LLMs in handling complex low-level tasks. On the other hand, we harness the high-level reasoning capabilities of LLMs to enhance low-level action performance within RL model training.
2.3 Large Language Models for Multi-Agent Systems
LLM-based Multi-Agent (LLM-MA) systems focus on diverse agent profiles, interactions, and collective decision-making. While this allows agents to collaborate on dynamic, complex tasks, it also increases computational overhead due to the communication between LLMs (Guo et al., 2024), (Sun et al., 2024). Camel Li et al. (2024) and MetaGPT Hong et al. (2023) employ multiple LLM agents to accomplish tasks like brainstorming and software development. Nascimento et al. (2023) enhance communication and agent autonomy by integrating GPT-based technologies. In multi-robot contexts, Chen et al. (2023) compare task success rates and token efficiencies of four multi-agent communication frameworks. SMART-LLM (Kannan et al., 2023) decompose multi-robot task plans into subgoals for LLM to enable efficient execution, while Co-NavGPT (Yu et al., 2023) integrates LLMs as global planners for cooperative navigation. Focusing on multi-agent pathfinding (MAPF), Chen et al. (2024) studies the performance of solving MAPF with LLMs. Agashe et al. (2023) proposes the LLM-Coordination (LLM-Co) to enable LLMs to play coordination games. Li et al. (2023) explore the use of LLMs in cooperative games within a text-based environment, and Ma et al. (2023) explores LLMs in the StarCraft II environment. In contrast, our method leverages the planning abilities of the LLM to train better MARL policies instead of using them directly as agents.
3 Problem Formulation
@@
Markov game (MG) is defined as a multi-agent decision-making problem when the interaction between multiple agents affect the state dynamics of the entire system and the reward of each agent under certain conditions (Littman, 1994). In this work, we consider a Markov game, or a stochastic game Owen (1982) defined as a tuple $G:=(\mathcal{N},S,A,\{r^{i}\}_{i\in\mathcal{N}},p,\gamma)$, where $\mathcal{N}$ is a set of $N$ agents, $S=S^{1}\times\cdots\times S^{N}$ is the joint state space, $A=A^{1}\times\cdots\times A^{N}$ is the joint action space, with $(S^{i},A^{i})$ as the state space and action space of agent $i$, respectively, $\gamma\in[0,1)$ is the discounting factor (Littman, 1994; Owen, 1982). The state transition $p:S\times A\rightarrow\Delta(S)$ is controlled by the current state and joint action, where $\Delta(S)$ represents the set of all probability distributions over the joint state space $S$. Each agent has a reward function, $r^{i}:S\times A\rightarrow\mathbb{R}$. At time $t$, agent $i$ chooses its action $a^{i}_{t}$ according to a policy $\pi^{i}:S\rightarrow\Delta(A^{i})$. For each agent $i$, it attempts to maximize its expected sum of discounted rewards, i.e. its objective function $J^{i}(s,\pi)=\mathbb{E}\left[\sum_{t=1}^{\infty}\gamma^{t-1}r_{t}^{i}(s_{t},a_{t})|s_{1}=s,a_{t}\sim\pi(\cdot|{s}_{t})\right]$. In the literature, deep MARL algorithms (Lowe et al., 2020; Yu et al., 2022; Rashid et al., 2018) have been designed to train neural network-based policies $\pi_{i}(\theta_{i})$. For a cooperative game, one shared reward function for all the agents is widely used during the training process, which is also considered in this work.
4 Methodology
@@
In this section, we introduce our method, YOLO-MARL, which leverages LLMs to enhance MARL. Specifically, during training, we utilize the high-level task planning capabilities of LLMs to guide the MARL process. Our approach consists of four key components: Strategy Generation, State Interpretation, Planning Function Generation, and MARL training process with the LLM generated Planning Function incorporated throughout.
@@
1:Large Language Model $LLM$, observation process function $F_{\text{S}}$, MARL actor $\mathcal{A}$, MARL algorithm $MARL_{alg}$, Prompts $P$
@@
3:// Sample functions code from the LLM
4:$\mathcal{F_{\mathcal{T}}}\sim LLM(P)$
5:for each training step do
6:     // Get processed global observation $S_{I}$ from $F_{\text{S}}$
7:     $S_{I}\leftarrow F_{\text{S}}(S_{v})$
8:     // Assign tasks $\mathcal{T}$ to each agent
9:     $\mathcal{T}_{1},\mathcal{T}_{2},\dotsc\leftarrow\mathcal{F_{\mathcal{T}}}(S_{I})$
10:     // Output actions from the actor
11:     $a_{1},a_{2},\dotsc\leftarrow\mathcal{A}(S_{v})$
12:     for each agent $i$ do
13:         if $a_{i}\in\mathcal{T}_{i}$ then
14:              $\Delta r_{i}\leftarrow r^{\prime}$
15:         else
16:              $\Delta r_{i}\leftarrow p^{\prime}$
17:         end if
18:     end for
19:     // Compute final reward for critic
20:     $R\leftarrow r+\sum_{i}\Delta r_{i}$
21:     // Use $R$ as the final reward for MARL training
22:     $\pi(\theta)=MARL_{alg}(R)$
23:end for
24:return Trained MARL policy
4.1 Strategy Generation
To create a generalizable framework applicable to various environments—especially when users may have limited prior knowledge—we incorporate a Strategy Generation Module into our methodology. This module enables the LLM to autonomously generate strategies for different environments without requiring extensive human input or expertise.
As shown inside the blue box of Figure 1(a), the LLM is provided with the basic information about the environment, including task descriptions, relevant rules, and constraints of how to interact with the environment. Additionally, we supply a general guideline within the prompt to assist the LLM in generating effective strategies. Gathering all the information, the LLM will output detailed strategies to accomplish the tasks or achieve the goal, following the specified format.
By aggregating all this information, the LLM outputs detailed strategies to accomplish the tasks or achieve the goals, following a specified format. The Strategy Generation is crucial for several reasons:
@@ 4.1 Strategy Generation
  • Facilitating Planning Function Generation: The strategies serve as vital components in the prompts used for the Planning Function Generation Module. The results of using YOLO-MARL but without Strategy Generation Module are shown in ablation study  6.1 .
The LLM-generated strategies are incorporated into the prompt alongside other necessary information to facilitate the subsequent planning function generation. Further details about the strategy prompts and their formats can be found in Appendix C.1.
4.2 State Interpretation
In many simulation environments, observations or states are typically provided as vectors, with each component constructed using various encoding methods. While the vector form of observation is easy to handle when training deep reinforcement learning models, it is difficult for LLMs to directly parse their semantic meaning due to the lack of explicit context for each component.
We propose the State Interpretation Module to assist the LLM in interpreting the environment state. By providing a semantically meaningful representation of the state, the LLM can successfully generate executable planning functions for training. Formally, given the current environment state in vector form $S_{v}$, we define an interpretation function $F_{S}$ such that $F_{S}(S_{v})\to S_{I}$, where $S_{I}$ provides more explicit and meaningful information about each state component. Specifically, we implement a Python function that processes the vector observation and restructures each component. For example, it extracts position information from the raw observation and organizes it into a variable name ”position”. This transformation enables the LLM to understand the context and significance of each part of the state.
Recent works like Ma et al. (2024) and Xie et al. (2023) have demonstrated the success of enhancing LLMs performance by providing relevant environment code. In the same manner, we include the interpretation function $F_{S}$ in the prompting pipeline, formatted as Pythonic environment code as shown in the purple box in Figure 1(b). The State Interpretation Module significantly reduces the risk of the LLM generating erroneous functions with outputs incompatible with the training procedures. An ablation study on the effectiveness of this module can be found in Sec 6.2, while more details about the interpretation function are provided in Appendix C.2.
4.3 Planning Function Generation

### v2

@@ Abstract
Advancements in deep multi-agent reinforcement learning (MARL) have positioned it as a promising approach for decision-making in cooperative games. However, it still remains challenging for MARL agents to learn cooperative strategies for some game environments. Recently, large language models (LLMs) have demonstrated emergent reasoning capabilities, making them promising candidates for enhancing coordination among the agents. However, due to the model size of LLMs, it can be expensive to frequently infer LLMs for actions that agents can take. In this work, we propose You Only LLM Once for MARL (YOLO-MARL), a novel framework that leverages the high-level task planning capabilities of LLMs to improve the policy learning process of multi-agents in cooperative games. Notably, for each game environment, YOLO-MARL only requires one time interaction with LLMs in the proposed strategy generation, state interpretation and planning function generation modules, before the MARL policy training process. This avoids the ongoing costs and computational time associated with frequent LLMs API calls during training. Moreover, trained decentralized policies based on normal-sized neural networks operate independently of the LLM. We evaluate our method across two different environments and demonstrate that YOLO-MARL outperforms traditional MARL algorithms. The Github repository of our code can be found at https://github.com/paulzyzy/YOLO-MARL.
I INTRODUCTION
@@
MARL has become a vital framework for solving decision-making problems in complex multi-agent systems. With the rising applications of multi-agent systems, it is increasingly crucial for individual agents to cooperate or compete in dynamic environments without relying on centralized control[1]. In cooperative Markov games, agents work together to maximize joint rewards. However, existing MARL approaches often struggle with learning effective distributed policies, particularly in tasks characterized by sparse rewards, dynamic environments, and large action spaces.
Concurrently, LLMs have demonstrated remarkable capabilities as high-level semantic planners by leveraging in-context learning and extensive prior knowledge[2]. Recent studies have showcased LLMs deployed as embodied agents [3, 4], as well as their use in guiding reinforcement learning (RL) as ELLM, which leverages LLMs to suggest goals [5], and work focusing on aligning LLM-provided actions with RL policies [6]. Despite these promising developments, integrating LLMs into multi-agent settings remains largely unexplored. Moreover, repeated interactions with LLMs during long-episode or complex tasks can be both time-consuming and computationally prohibitive, particularly when training over tens of millions of steps.
@@
Addressing these challenges, we propose YOLO-MARL, a novel framework combining LLM’s high-level planning capabilities with MARL policy training, as shown in Fig. 1. YOLO-MARL uniquely requires just one LLM interaction per environment. Once the strategy generation, state interpretation, and planning function generation modules produce the necessary guidance, the MARL training process proceeds without further LLM involvement. This design substantially reduces the communication overhead and computational cost typically associated with LLM inferences. Moreover, YOLO-MARL demonstrates its strong generalization capability and simplicity for application, since it only requires a basic understanding of the new environments from users. We validate our framework on a sparse reward multi-agent environment: Level-Based Foraging (LBF) environment[1] as well as the Multi-Agent Particle (MPE) environment[7], and our results show that YOLO-MARL outperforms several MARL baselines, such as[8, 9, 7]. To the best of our knowledge, this work represents one of the first trials to fuse the high-level reasoning and planning abilities of LLMs with MARL, pointing a new direction for scalable and efficient multi-agent coordination [10].
@@
  • This framework synergizes the planning capabilities of LLMs with MARL to enhance the policy learning in challenging cooperative game environments. In particular, our approach exploits the LLM’s wide-ranging reasoning ability to generate high-level planning functions to facilitate agents in coordination.
@@
An overview of YOLO-MARL is presented in Fig. 1, and all related prompts, environments, and generated planning functions are available in our GitHub repository.
@@
Figure 1: Depiction of our framework YOLO-MARL. (a). Strategy Generation: We pass basic environment and task description into the LLM to get generated strategies for this specific environment. (b). State Interpretation: We process the global states so that the format of global states will be more structured and organized for better comprehension by the LLM. (c). Planning Function Generation: We chain together the environment and task description, LLM generated strategies and state interpretation function. These prompts are then fed into the LLM to generate a Python planning function for this environment. (d). MARL Training: The state interpretation function and the generated planning function are integrated into the MARL training process. The LLM is no longer required for further interaction after the Planning Function Generation. The more detailed explanation of MARL Training part can be found in Algorithm  1
II Related Work
II-A Multi-Agent Reinforcement Learning
MARL has attracted significant attention for its potential to address complex, decentralized problems. A popular paradigm is centralized training with decentralized execution. Methods like QMIX[9] and MADDPG[7] employ centralized critics during training to coordinate agents, while allowing independent execution at test time. In cooperative environments, COMA[11] and VDN[12] enable agents to share rewards and maximize joint returns. Despite these advances, many existing MARL algorithms struggle in sparse-reward settings and have difficulty learning fully cooperative policies. Moreover, only limited work has explored the integration of LLMs with MARL[10], leaving open questions about leveraging LLMs for multi-agent decision-making.
II-B Large Language Models for RL and Decision-Making
Recent studies have incorporated LLMs into the RL training process to enhance performance. For instance, [5] improve agent exploration by aligning LLM-suggested goals with observed behaviors, while [13] generate actions conditioned on language-based prompts during online RL. Similarly, [6] uses LLMs to provide scalar rewards that guide training. However, many of these approaches focus on single-agent settings or require extensive interactions with LLMs during training. And [14] highlights the limitations of LLMs in handling complex low-level tasks. Other works, such as [15], studies a multi-task RL problem. SayCan [2] grounds LLMs via value functions of pretrained skills to execute abstract commands on robots. [16] finds that code-writing LLMs can be re-purposed to write robot policy code. Additionally, studies like [17] and [18] exploit LLMs’ prior knowledge and code-generation capabilities to produce reward functions. In contrast, our approach leverages LLMs to generate planning functions, thereby enhancing MARL without continuous LLM involvement.
II-C Large Language Models for Multi-Agent Systems
LLM-based multi-agent systems have been employed in diverse applications requiring varied agent roles and collaborative decision-making[19, 10]. Camel[20] and MetaGPT[21] deploy multiple LLM agents for tasks like brainstorming and software development. SMART-LLM[22] decomposes multi-robot task planning into subgoals, and Co-NavGPT[23] uses LLMs as global planners for cooperative navigation. Other research has focused on applying LLMs in strategic gaming environments[24, 25, 26, 27, 28]. Unlike these approaches, which use LLMs directly as agents or decision-makers, our method harnesses the planning capabilities of LLMs to train compact, efficient MARL policies.
III Problem Formulation
@@
Markov game is defined as a multi-agent decision-making problem when the interaction between multiple agents affect the state dynamics of the entire system and the reward of each agent under certain conditions [29]. In this work, we consider a Markov game, or a stochastic game [30] defined as a tuple $G:=(\mathcal{N},S,A,\{r^{i}\}_{i\in\mathcal{N}},p,\gamma)$, where $\mathcal{N}$ is a set of $N$ agents, $S=S^{1}\times\cdots\times S^{N}$ is the joint state space, $A=A^{1}\times\cdots\times A^{N}$ is the joint action space, with $(S^{i},A^{i})$ as the state space and action space of agent $i$, respectively, $\gamma\in[0,1)$ is the discounting factor[29, 30]. The state transition $p:S\times A\rightarrow\Delta(S)$ is controlled by the current state and joint action, where $\Delta(S)$ represents the set of all probability distributions over the joint state space $S$. Each agent has a reward function, $r^{i}:S\times A\rightarrow\mathbb{R}$. At time $t$, agent $i$ chooses its action $a^{i}_{t}$ according to a policy $\pi^{i}:S\rightarrow\Delta(A^{i})$.
For each agent $i$, it attempts to maximize its expected sum of discounted rewards, i.e. its objective function $J^{i}(s,\pi)=\mathbb{E}\left[\sum_{t=1}^{\infty}\gamma^{t-1}r_{t}^{i}(s_{t},a_{t})|s_{1}=s,a_{t}\sim\pi(\cdot|{s}_{t})\right]$. In the literature, MARL algorithms[7, 8, 9] have been designed to train neural network-based policies $\pi_{i}(\theta_{i})$. For a cooperative game, one shared reward function for all agents is widely used during the training process, which is also considered in this work.
IV Methodology
@@
In this section, we introduce YOLO-MARL, a framework that leverages the high-level planning capabilities of LLMs to enhance MARL. YOLO-MARL integrates four key components: Strategy Generation, State Interpretation, Planning Function Generation, and MARL training process with the LLM generated Planning Function incorporated throughout.
@@
1:Large Language Model $LLM$, State Interpretation function $F_{\text{S}}$, MARL actor $\mathcal{A}$, MARL algorithm $MARL_{alg}$, Initial Prompts $P_{init}$
@@
3:$P_{Strategy}\sim LLM(P_{init})$ // Strategy Generation
4:$P=P_{init}+P_{Strategy}+F_{\text{S}}$ // Chaining all the prompt for Planning Function Generation
5:$\mathcal{F_{\mathcal{T}}}\sim LLM(P)$ // Planning Function Generation: Sample functions code from the LLM
6:MARL training with generated planning function
7:for each training step do
8:    $S_{I}\leftarrow F_{\text{S}}(S_{v})$ // State Interpretation: Get processed global observation $S_{I}$ from $F_{\text{S}}$
9:    $\mathcal{T}_{1},\mathcal{T}_{2},\dotsc\leftarrow\mathcal{F_{\mathcal{T}}}(S_{I})$ // Assign tasks $\mathcal{T}$ to each agent
10:    $a_{1},a_{2},\dotsc\leftarrow\mathcal{A}(S_{v})$ // Output actions from the actor
11:    for each agent $i$ do
12:         if $a_{i}\in\mathcal{T}_{i}$ then
13:             $\Delta r_{i}\leftarrow r^{\prime}$
14:         else
15:             $\Delta r_{i}\leftarrow p^{\prime}$
16:         end if
17:    end for
18:    $R\leftarrow r+\sum_{i}\Delta r_{i}$ // Compute final reward for critic: More details are in equation  12
19:    $\pi(\theta)=MARL_{alg}(R)$ // Use $R$ as the final reward for MARL training
20:end for
21:return Trained MARL policy
IV-A Strategy Generation
To ensure applicability across diverse environments—even for users with limited domain knowledge—we incorporate a Strategy Generation Module, as shown in the blue box of Figure 1(a). In this module, the LLM receives basic environment details (e.g., task descriptions, rules, and constraints) along with a general guideline, and then autonomously outputs a detailed strategy in a prescribed format without requiring extensive human input or expertise.
The Strategy Generation is crucial for several reasons:
@@ IV-A Strategy Generation
  • Facilitating Planning Function Generation: The strategies serve as vital components in the prompts used for the Planning Function Generation Module. The results of using YOLO-MARL but without Strategy Generation Module are shown in ablation study  VI-A .
The LLM-generated strategies are incorporated into the prompt alongside other necessary information to facilitate the subsequent planning function generation.
IV-B State Interpretation
In many simulation environments, observations are provided as vectors, with each component encoded in a non-semantic format. Although such representations are effective for training deep RL models, they pose challenges for LLMs that require context to interpret each component correctly.
We propose the State Interpretation Module to assist the LLM in interpreting the environment state. By providing a semantically meaningful representation of the state, the LLM can successfully generate executable planning functions for training. Formally, given the current environment state in vector form $S_{v}$, we define an interpretation function $F_{S}$ such that $F_{S}(S_{v})\to S_{I}$, where $S_{I}$ provides more explicit and meaningful information about each state component.
Recent works like [17] and [18] have demonstrated the success of enhancing LLMs performance by providing relevant environment code. In the same manner, we include the interpretation function $F_{S}$ in the prompting pipeline, formatted as Python environment code as shown in the purple box in Figure 1(b). The State Interpretation Module significantly reduces the risk of the LLM generating erroneous functions with outputs incompatible with the training procedures. An ablation study on the effectiveness of this module can be found in Sec VI-B.
IV-C Planning Function Generation